The perimeter of a rectangle is 22. The width of the rectangle is one less than the square of the length. Find the dimensions of the rectangle.
Answer by stanbon(75887)

The perimeter of a rectangle is 22.
The width of the rectangle is one less than the square of the length.
Let the length be "x".
Then width is "x^2-1"
-------
Find the dimensions of the rectangle.
Perimeter = 2(length + width)
22 = 2(x + x^2-1)
11 = x^2+x-1
x^2+x-12 = 0
(x+4)(x-3) = 0
Positive solution:
x = 3 (length)
x^2-1 = 9-1 = 8 (width)
===========================
cheers,
Stan H.